OpenStudy (samara89):

Zuri made 2 spheres using plaster. The smaller sphere has a radius of 6 cm. The larger sphere has a radius of 24 cm. How much more plaster did Zuri use for the larger sphere? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (samara89):

@welshfella

OpenStudy (samara89):

do you add or subtract the answers on this one

OpenStudy (greatlife44):

you need to subtract them

OpenStudy (greatlife44):

\[\frac{ 4 }{ 3 }\pi*r_{sphere~A}^{3} - \frac{ 4 }{ 3 }\pi*r_{sphere B}^{3}\] Sphere A r = 24 Sphere B r = 6

OpenStudy (samara89):

57876.48-904.32

OpenStudy (greatlife44):

correct make sure you keep your answer in cubic centimetes

OpenStudy (samara89):

and I get 56972.16

OpenStudy (samara89):

is that correct

OpenStudy (greatlife44):

Yeah but make sure you report your answer in centimeters cubed

OpenStudy (samara89):

can you help me with more

OpenStudy (greatlife44):

yeah, tag me